OpenStudy (anonymous):

What is the square root of 42

OpenStudy (anonymous):

You usually need a calculator. That should help.

OpenStudy (anonymous):

i got 6.480 what does other people think?

OpenStudy (anonymous):

That's right^^

OpenStudy (anonymous):

I go 4, 2, sqrt on my calc and it says 6.48blah

OpenStudy (anonymous):

Yeah its 6.48 there is no way to simplify this one like some other square root numbers

OpenStudy (anonymous):

\[x_{n+1} = \frac{21}{x_n} + \frac{x_n}{2}\]

OpenStudy (anonymous):

if ya dont have a calculator

OpenStudy (precal):

 notice that 2 and 3 and 7 are not perfect squares and that the square root of 42 is in its simplest form